OpenStudy (anonymous):

which expression is equal to the expression below ? (7*2)^7 a.7^7+2^2 b.9^7 c.7^*2^7 d.7*( 7*2) ???? help plz thank u : )

OpenStudy (anonymous):

c) if its 7^7*2^7

OpenStudy (anonymous):

think there is a typo in the answers

Directrix (directrix):

(7*2)^7 = 7^7 * 2^7 These will be quick and easy for you if you know the properties for exponential expressions. Product to a power: (ab)m = a^m*b^m To raise a product to a power, raise each factor to the power. Example: (x4 y5 )3 = x12 y15 http://www.rock.uwc.edu/academics/tutoring/lsc/handouts/Properties%20of%20Exponents.pdf